OpenStudy (cutiecomittee123):

f(x)=3/x-2 -sqrt x-3 f(19)= so i solved for F(19) and I got 1/7-4 but i tried typing that into this test I am taking and it said invalid format. Then I combined like terms and i got -27/7 whih was still in the wrong format, so now what?

OpenStudy (tkhunny):

Do you mean \(\dfrac{3}{x-2}\;or\;\dfrac{3}{x}-2\)? You have written the latter.

OpenStudy (anonymous):

3/(x-2-sqrt(x-3) is this the equation? i get 3/(17-4) =3/13

OpenStudy (cutiecomittee123):

no the question is 3 divided by x-2 minus the quare root of x-3

OpenStudy (anonymous):

try decimal form

OpenStudy (anonymous):

-3.8 is what i got

OpenStudy (cutiecomittee123):

Alright thanks